Year Nominal GDP(Billions of current dollars Real GDP(Billions of dollars of 2009) 2013 16,691.50 15,612.20 2014...

Year

Nominal

GDP(Billions of current dollars

Real

GDP(Billions of dollars of 2009)

2013

16,691.50 15,612.20
2014 17,393.10 15,982.30
2015 18,036.60 16,397.20
2016 18,569.10 16,662.10

Compute the rate of growth of the economy (use real GDP), and briefly comment on your results.

-Between 2013 and 2014

-Between 2014 and 2015

-Between 2015 and 2016

Which one is more important: Nominal GDP or Real GDP? Explain Why.
Answer #1

1> Between 2013 and 2014. Growth rate is (Real GDP in 2014 - Real GDP in 2013) / Real GDP in 2013

= (17,393.10-16,691.50)/16,691.50

=4.203337%

2> Between 2014 and 2015. Growth rate is (Real GDP in 2015 - Real GDP in 2014) / Real GDP in 2014

= (18,036.60-17,393.10)/17,393.10

=3.699743%

3> Between 2015 and 2016. Growth rate is (Real GDP in 2016 - Real GDP in 2015) / Real GDP in 2015

= (18,569.10-18,036.60)/18,036.60

=2.95233%

Comment - We can see, that the growth rate has fallen over the years. So, the economy is growing slower than before.

4> Real GDP is more important because it does not consider the growth in GDP due to the change in price, so the effect of price on GDP is eliminated, so real GDP is more important.